# The algebraic and geometric classification of

### v1

@@ Introduction
The algebraic classification (up to isomorphism) of algebras of small dimensions from a certain variety defined by a family of polynomial identities is a classic problem in the theory of non-associative algebras. Another interesting approach to studying algebras of a fixed dimension is to study them from a geometric point of view (that is, to study the degenerations and deformations of these algebras). The results in which the complete information about degenerations of a certain variety is obtained are generally referred to as the geometric classification of the algebras of these varieties. There are many results related to the algebraic and geometric classification of Jordan, Lie, Leibniz, Zinbiel, and other algebras (see, [3, 20, 8, 9, 12, 13, 14, 16, 15, 1] and references in [17, 25, 18]).
@@
The class of noncommutative Jordan algebras turned out to be vast: for example, apart from alternative and Jordan algebras, it contains quasi-associative and quasi-alternative algebras, quadratic flexible algebras and all anticommutative algebras. However, the structure theory of this class is far from being nice. Nevertheless, certain progress was made in the study of the structure theory of noncommutative Jordan algebras and superalgebras (see, for example [5, 35, 26, 27, 19, 30, 4, 23, 24, 28, 31, 32, 33, 39, 29]). So, Schafer gave the first structure theory for noncommutative Jordan algebras of characteristic zero in [35]. Kleinfeld and Kokoris proved that a simple, flexible, power-associative algebra of finite dimension over a field of characteristic zero is a noncommutative Jordan algebra [22]. A coordinatization theorem for noncommutative Jordan algebras was obtained in a paper by McCrimmon [26]. Later, McCrimmon shows that the property of being a noncommutative Jordan algebra is preserved under homotopy [27]. Noncommutative Jordan algebras with additional identity $([x,y],y,y)=0$ were studied in papers by Shestakov and Schafer [39, 36]. Strongly prime noncommutative Jordan algebras were studied by Skosyrskiĭ [37, 38]. A connection between noncommutative Jordan algebras and $(-1,-1)$-balanced Freudenthal Kantor triple systems was established by Elduque, Kamiya, and Okubo [10]. Cabrera Serrano, Cabrera García, and Rodríguez Palacios studied the algebra of multiplications of prime and semiprime noncommutative Jordan algebras [4, 5]. Jumaniyozov, Kaygorodov, and Khudoyberdiyev classified complex $4$-dimensional nilpotent noncommutative Jordan algebras [16]. Connections between noncommutative Jordan algebras and conservative algebras are discussed in [29]. There are also some results in noncommutative Jordan superalgebras. Namely, Pozhidaev and Shestakov classified all simple finite-dimensional noncommutative Jordan superalgebras in [32, 33]; Kaygorodov, Popov, and Lopatin studied the structure of simple noncommutative Jordan superalgebras [19]; Popov described representations of simple noncommutative Jordan superalgebras in [30, 31].
@@
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$. It is defined by $5$ rigid algebras and three one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 23. The variety of complex $4$-dimensional nilpotent noncommutative Jordan algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [16, Theorem B].
@@
The variety of complex $3$-dimensional standard algebras has dimension $9$. It is defined by $13$ rigid algebras and one one-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 22. The variety of complex $4$-dimensional nilpotent standard algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [16, Theorem B].
@@ 1.3. The algebraic classification of Kokoris algebras
A flexible algebra $\bf A$ is called a Kokoris algebra if $\bf A^{+}$ is associative. Kokoris algebras were introduced in [24] and they also appear in [40, 38, 7]. $({\bf A},\cdot)$ is a Kokoris algebra if and only if $({\bf A},\circ,[\cdot,\cdot])$ is a generic Poisson algebra. Also, observe that Poisson algebras are related to Lie admissible Kokoris algebras.
@@ 2.3. The geometric classification of standard algebras
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{16},$ ${\bf A}_{19}^{0}\}$$\not\to$ ${\bf A}_{19}^{\frac{1}{2}},$ ${\bf A}_{19}^{-\frac{1}{2}},$$\{{\bf A}_{17}^{\frac{1}{2}},$ ${\bf A}_{14}^{0,\frac{1}{2}},$ ${\bf A}_{14}^{0,-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},$ ${\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{02}^{\alpha}\}.$
@@ 2.4. The geometric classification of noncommutative Jordan algebras
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$ and it has $8$ irreducible components defined by
@@
$\mathcal{C}_{5}=\overline{\mathcal{O}({\bf A}_{19}^{\alpha})},$  $\mathcal{C}_{6}=\overline{\mathcal{O}({\bf A}_{24}^{\alpha})},$  $\mathcal{C}_{7}=\overline{\mathcal{O}({\bf A}_{30})},$  $\mathcal{C}_{8}=\overline{\mathcal{O}({\bf A}_{32})}.$
In particular, there are only five rigid algebras in this variety.
@@
${\bf A}_{04}$ and ${\bf A}_{12}$ are commutative, ${\bf A}_{24}^{\alpha}$ are anticommutative, ${\bf A}_{32},$ ${\bf A}_{30},$ ${\bf A}_{19}^{\alpha}$ and ${\bf A}_{17}^{\alpha}$ are noncommutative and non-anticommutative. Hence,
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{24}^{\alpha}\}$$\not\to$ $\{{\bf A}_{32},$ ${\bf A}_{30},$ ${\bf A}_{19}^{\alpha},$ ${\bf A}_{17}^{\alpha}\}.$
@@
${\bf A}_{17}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16},{\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\ A_{3}A_{3}=0,\\ c_{22}^{3}=0,\ c_{21}^{2}=c_{12}^{2},\ c_{12}^{2}c_{23}^{3}+c_{11}^{1}c_{32}^{3}=c_{13}^{3}(c_{23}^{3}+c_{32}^{3}),\\ c_{23}^{3}+c_{32}^{3}=c_{22}^{2},\ c_{21}^{3}c_{32}^{3}=c_{12}^{3}c_{23}^{3},\ c_{12}^{2}c_{21}^{3}=c_{11}^{3}c_{23}^{3}\par\par\end{array}\right\}$
${\bf A}_{19}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16},{\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\\ \ A_{2}A_{3}+A_{3}A_{2}=0,\ c_{31}^{3}+c_{13}^{3}=c_{11}^{1}\end{array}\right\}$
${\bf A}_{32}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{12},\ {\bf A}_{16},\ {\bf A}_{17}^{\alpha},\\ {\bf A}_{19}^{\alpha},\ {\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}c_{11}^{2}=0,\ c_{11}^{3}=0,\ c_{22}^{1}=0,\ c_{22}^{3}=0,\ c_{33}^{1}=0,\ c_{33}^{2}=0,\\ c_{12}^{3}=-c_{21}^{3},\ c_{21}^{1}+c_{12}^{1}=c_{22}^{2},\ c_{21}^{2}+c_{12}^{2}=c_{11}^{1},\\ c_{23}^{3}+c_{32}^{3}=c_{22}^{2},\ c_{31}^{3}+c_{13}^{3}=c_{11}^{1}\par\par\par\par\end{array}\right\}$
@@
    ${\bf A}_{19}^{\frac{\alpha}{2}}\xrightarrow{(te_{1},-2t^{2}e_{2}+e_{3},t^{3}e_{2})}{\bf A}_{02}^{\alpha},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1}+e_{2},e_{3},te_{2})}{\bf A}_{13}^{\alpha},$  ${\bf A}_{33}^{\beta}\xrightarrow{(e_{1}+(\beta-\alpha)e_{2}-\alpha e_{3},t^{2}e_{3},te_{2}+te_{3})}{\bf A}_{14}^{\alpha,\beta},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1},e_{3},te_{2})}{\bf A}_{18}^{\alpha},$  ${\bf A}_{17}^{\frac{-2-t}{2t}}\xrightarrow{(e_{1}+e_{2},te_{2},e_{3})}{\bf A}_{29},$  ${\bf A}_{33}^{\frac{2\alpha-t}{2+2t}}\xrightarrow{((1+t)e_{1}-(\alpha+\frac{t}{2})e_{3},-t^{2}e_{3},e_{2}+te_{3})}{\bf A}_{31}^{\alpha},$  ${\bf A}_{32}\xrightarrow{(e_{1}-\alpha e_{3},te_{3},(\alpha+t^{2})e_{2}+e_{3})}{\bf A}_{33}^{\alpha}.$

### v2

@@ Introduction
The algebraic classification (up to isomorphism) of algebras of small dimensions from a certain variety defined by a family of polynomial identities is a classic problem in the theory of non-associative algebras. Another interesting approach to studying algebras of a fixed dimension is to study them from a geometric point of view (that is, to study the degenerations and deformations of these algebras). The results in which the complete information about degenerations of a certain variety is obtained are generally referred to as the geometric classification of the algebras of these varieties. There are many results related to the algebraic and geometric classification of Jordan, Lie, Leibniz, Zinbiel, and other algebras (see, [4, 21, 9, 10, 13, 14, 15, 17, 16, 1] and references in [18, 26, 19]). The geometric classification of algebras from a certain variety is based on the notion of degeneration, that is a ”dual” notion to deformations [3, 29].
@@
The class of noncommutative Jordan algebras turned out to be vast: for example, apart from alternative and Jordan algebras, it contains quasi-associative and quasi-alternative algebras, quadratic flexible algebras and all anticommutative algebras. However, the structure theory of this class is far from being nice. Nevertheless, certain progress was made in the study of the structure theory of noncommutative Jordan algebras and superalgebras (see, for example [6, 36, 27, 28, 20, 32, 5, 24, 25, 30, 33, 34, 40, 31]). So, Schafer gave the first structure theory for noncommutative Jordan algebras of characteristic zero in [36]. Kleinfeld and Kokoris proved that a simple, flexible, power-associative algebra of finite dimension over a field of characteristic zero is a noncommutative Jordan algebra [23]. A coordinatization theorem for noncommutative Jordan algebras was obtained in a paper by McCrimmon [27]. Later, McCrimmon shows that the property of being a noncommutative Jordan algebra is preserved under homotopy [28]. Noncommutative Jordan algebras with additional identity $([x,y],y,y)=0$ were studied in papers by Shestakov and Schafer [40, 37]. Strongly prime noncommutative Jordan algebras were studied by Skosyrskiĭ [38, 39]. A connection between noncommutative Jordan algebras and $(-1,-1)$-balanced Freudenthal Kantor triple systems was established by Elduque, Kamiya, and Okubo [11]. Cabrera Serrano, Cabrera García, and Rodríguez Palacios studied the algebra of multiplications of prime and semiprime noncommutative Jordan algebras [5, 6]. Jumaniyozov, Kaygorodov, and Khudoyberdiyev classified complex $4$-dimensional nilpotent noncommutative Jordan algebras [17]. Connections between noncommutative Jordan algebras and conservative algebras are discussed in [31]. There are also some results in noncommutative Jordan superalgebras. Namely, Pozhidaev and Shestakov classified all simple finite-dimensional noncommutative Jordan superalgebras in [34]; Kaygorodov, Popov, and Lopatin studied the structure of simple noncommutative Jordan superalgebras [20]; Popov described representations of simple noncommutative Jordan superalgebras in [32, 33].
@@
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$. It is defined by $4$ rigid algebras and three one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 23. The variety of complex $4$-dimensional nilpotent noncommutative Jordan algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [17, Theorem B].
@@
The variety of complex $3$-dimensional standard algebras has dimension $9$. It is defined by $13$ rigid algebras and one one-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 22. The variety of complex $4$-dimensional nilpotent standard algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [17, Theorem B].
@@ 1.3. The algebraic classification of Kokoris algebras
A flexible algebra $\bf A$ is called a Kokoris algebra if $\bf A^{+}$ is associative. Kokoris algebras were introduced in [25] and they also appear in [39, 8]. $({\bf A},\cdot)$ is a Kokoris algebra if and only if $({\bf A},\circ,[\cdot,\cdot])$ is a generic Poisson algebra. Also, observe that Poisson algebras are related to Lie admissible Kokoris algebras.
@@ 2.3. The geometric classification of standard algebras
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{16},$ ${\bf A}_{19}^{0}\}$$\not\to$ $\big{\{}{\bf A}_{19}^{\frac{1}{2}},$ ${\bf A}_{19}^{-\frac{1}{2}},$ ${\bf A}_{17}^{\frac{1}{2}},$ ${\bf A}_{14}^{0,\frac{1}{2}},$ ${\bf A}_{14}^{0,-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},$ ${\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{02}^{\alpha}\big{\}}.$
@@ 2.4. The geometric classification of noncommutative Jordan algebras
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$ and it has $7$ irreducible components defined by
@@
$\mathcal{C}_{5}=\overline{\mathcal{O}({\bf A}_{19}^{\alpha})},$  $\mathcal{C}_{6}=\overline{\mathcal{O}({\bf A}_{24}^{\alpha})},$  $\mathcal{C}_{7}=\overline{\mathcal{O}({\bf A}_{32})}.$
In particular, there are only four rigid algebras in this variety.
@@
${\bf A}_{04}$ and ${\bf A}_{12}$ are commutative, ${\bf A}_{24}^{\alpha}$ are anticommutative, ${\bf A}_{32},$ ${\bf A}_{19}^{\alpha}$ and ${\bf A}_{17}^{\alpha}$ are noncommutative and non-anticommutative. Hence,
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{24}^{\alpha}\}$$\not\to$ $\{{\bf A}_{32},$ ${\bf A}_{19}^{\alpha},$ ${\bf A}_{17}^{\alpha}\}.$
@@
${\bf A}_{17}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{2}^{2}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\ c_{22}^{3}=0,\ c_{23}^{3}+c_{32}^{3}=c_{22}^{2}\par\par\end{array}\right\}$
${\bf A}_{19}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\\ \ A_{2}A_{3}+A_{3}A_{2}=0,\ c_{31}^{3}+c_{13}^{3}=c_{11}^{1}\end{array}\right\}$
${\bf A}_{32}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{12},\ {\bf A}_{16},\\ {\bf A}_{17}^{\alpha},\ {\bf A}_{19}^{\alpha}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{2}^{2}\subset A_{2},\ c_{23}^{2}=-c_{32}^{2},\ c_{23}^{3}=-c_{32}^{3},\ c_{22}^{2}=0,\ c_{22}^{3}=0,\ c_{33}^{2}=0,\ c_{33}^{3}=0\par\par\par\end{array}\right\}$
@@
    ${\bf A}_{19}^{\frac{\alpha}{2}}\xrightarrow{(te_{1},-2t^{2}e_{2}+e_{3},t^{3}e_{2})}{\bf A}_{02}^{\alpha},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1}+e_{2},e_{3},te_{2})}{\bf A}_{13}^{\alpha},$
    ${\bf A}_{33}^{\beta}\xrightarrow{(e_{1}+(\beta-\alpha)e_{2}-\alpha e_{3},t^{2}e_{3},te_{2}+te_{3})}{\bf A}_{14}^{\alpha,\beta},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1},e_{3},te_{2})}{\bf A}_{18}^{\alpha},$  ${\bf A}_{17}^{\frac{-2-t}{2t}}\xrightarrow{(e_{1}+e_{2},te_{2},e_{3})}{\bf A}_{29},$  ${\bf A}_{19}^{-\frac{1}{2}+\frac{1}{t}}\xrightarrow{(e_{2},te_{1},e_{3})}{\bf A}_{30},$  ${\bf A}_{33}^{\frac{2\alpha-t}{2+2t}}\xrightarrow{((1+t)e_{1}-(\alpha+\frac{t}{2})e_{3},-t^{2}e_{3},e_{2}+te_{3})}{\bf A}_{31}^{\alpha},$  ${\bf A}_{32}\xrightarrow{(e_{1}-\alpha e_{3},te_{3},(\alpha+t^{2})e_{2}+e_{3})}{\bf A}_{33}^{\alpha}.$